# AIMCoT: Active Information-driven Multimodal Chain-of-Thought for Vision-Language Reasoning

### v1

@@ 4.1 Preliminaries
Vision-Language Model. A VLM typically fuses a vision encoder for preprocessing visual input and a generative language model, which jointly enable it to respond in a human-like manner as follows:
$answer=\text{VLM}(I,x),$ (1)
where $I$ and $x$ are the image and query, respectively.
Multimodal CoT. Compared to the direct response shown in Equation 1, multimodal CoT encourages the VLM to output the thought process, called rationales, before outputting the final answer. Existing work Gao et al. (2025) has been devoted to generalizing rationales from pure text to text-vision interleaved form.
@@ 4.2 Context-enhanced Attention-map Generation (CAG)
In this component, the VLM is prompted to carefully generate an explanatory description of the given image within the question before the process of VQA, with the explicit goal of helping a potential respondent correctly answer the question. In this process, the VLM acts as a facilitator who interprets the image in the context of the given question to guide the respondent’s thought process. Formally, it is expressed as follows:
$\mathcal{D}_{CAG}=\text{VLM}(I,x,\mathcal{P}_{CAG}),$ (2)
where $\text{VLM},I,x$ are the used VLM, the given image and question, respectively. $\mathcal{P}_{CAG}$ is the prompt provided for the model to generate the description. Then, to compensate for the sparsity of textual information within the context, the generated description $\mathcal{D}_{CAG}$ is concatenated to the question $x$ as follows:
$x^{\prime}=\text{concat}(x,\mathcal{D}_{CAG}).$ (3)
By enhancing the context to compensate for the sparsity of textual information, the disparity in text-vision granularity is effectively mitigated, unlocking the potential for the attention map to serve as a more reliable indicator of task-relevant regions. We provide an example in Appendix G.1 that details the template of $\mathcal{P}_{CAG}$, the entire process of CAG, and how the final attention map $A^{\prime}$ is generated, which is used in the next stage.
@@ 4.3 Active Visual Probing (AVP)
Built upon the theoretical foundation of information gain, AVP is designed to select the crucial regions from a set of candidate visual regions. Although AVP consists of three steps elaborated upon as follows, in terms of complexity, our provided analysis and empirical results in Appendices J and L suggest that the introduction of AVP still enables AIMCoT to strike a good balance between deployability and superior performance. A visualization of AVP is shown in Figure 3.
Figure 3: An overview of our AVP module, which iteratively selects $K$ most informative regions from a diversified candidate set $C$ to build an interleaved CoT that fosters vision-language reasoning.
Diversified Set of Candidate Regions Construction. Based on the first two insights summarized in Section 4, an inference can be obtained: relying solely on attention maps as the source may prevent the optimal region selection. Motivated by this, we propose to diversify the source of the candidate regions, thereby reducing the model’s dependency on attention maps alone. Specifically, we not only construct an attention-driven candidate set, $C_{attn}$, by selecting $N$ regions with the highest attention scores ($N\in\mathds{R}$), but also generate an exploratory candidate set, $C_{exp}$, by sampling $M$ grid regions uniformly at random from the input image ($M\in\mathds{R}$). Our empirical analysis in Appendix I shows that the incorporation of exploratory set $C_{exp}$ provides substantial salient visual regions for the VLM. Furthermore, we study the impact of introducing $C_{exp}$ on the performance of AIMCoT, and further compare random sampling, selective search Uijlings et al. (2013), and FastSAM Zhao et al. (2023) as methods for constructing $C_{exp}$ in Appendix M.
Formally, the process can be expressed as follows:
$C_{attn}$ $=\{R_{1},R_{2},\cdots,R_{N}\},~~~\text{s.t. $R_{i}$=Top-$i$ Region from the attention map $A^{\prime}$},$ (4)
$C_{exp}$ $=\{R_{N+1},R_{N+2},\cdots,R_{N+M}\},~~~~~C=C_{attn}\cup C_{exp},$ (5)
where $C_{attn},C_{exp}$, and $C$ are the attention-driven, exploratory, and total candidate sets, respectively. Subsequently, the goal is to select the most informative $K$ regions from $C$ ($K<|C|=N+M$).
Quantification of Information Gain. An ideal region of input image is the one that is most informative and critical for answering the question, i.e., the one that minimizes the model’s predictive uncertainty. Based on this idea, the region selection problem is transformed into a sequential information gain maximization task. To formalize this objective, we propose the concept of basic uncertainty $U_{B}$ defined as follows:
$U_{B}=H(Y|I,x,y_{<t})=-\sum_{y\in V}P(y|I,x,y_{<t})\text{log}_{2}P(y|I,x,y_{<t}),$ (6)
where $I,x$ and $y_{<t}$ are the input image, question and the predicted tokens, respectively. $P(y|I,x,y_{<t})$ is model’s probability distribution across the vocabulary $V$ in current context without the introduction of any region $R_{i}\in C$. It is seen that $U_{B}$ takes the form of entropy, thereby inherently capturing the model’s uncertainty when predicting the next token $t$ without incorporating any $R_{i}\in C$.
Similarly, the conditional uncertainty $U_{C,i}$, which illustrates the model’s uncertainty when an arbitrary region $R_{i}\in C$ is included into the context, is proposed and defined as follows:
$U_{C,i}=H(Y|I,x,y_{<t},R_{i})=-\sum_{y\in V}P(y|I,x,y_{<t},R_{i})\text{log}_{2}P(y|I,x,y_{<t},R_{i}),$ (7)
where $P(y|I,x,y_{<t},R_{i})$ is the model’s probability distribution across the vocabulary $V$ in the context combined with the region $R_{i}$. Eventually, the aforementioned conceptual formulation logically motivates the definition of our final key metric: the information gain of region $R_{i}$, which is formally defined as follows:
$IG(\{R_{i}\})=U_{B}-U_{C,i},~~~i=1,2,\cdots,N+M.$ (8)
Intuitively, $IG(\{R_{i}\})$ quantitatively characterizes how incorporating region $R_{i}$ reduces the model’s uncertainty for the subsequent token prediction.
Optimal Region Selection. To achieve the target of region selection, we formalize the problem as follows:
$\text{Maximize}~~F(S)=IG(S),~~~\text{s.t.}~S\subset C,|S|=K,$ (9)
where $S$ is the optimal set of selected regions and $K$ is the size of $S$ serving as a hyper-parameter. Motivated by two crucial insights, we propose a greedy algorithm detailed in Algorithm 1 to solve this problem: (1) According to recognized works Bian et al. (2017); Sener & Savarese (2018); Kim et al. (2016); Krause et al. (2008); Das & Kempe (2011), the greedy algorithm is a well-established and widely-used method for maximizing functions, especially for maximizing those that exhibit a tendency towards submodularity, even if they are not submodular theoretically; (2) Our experiments in Appendix H suggest that the information gain function $F$ empirically exhibits significant approximate submodularity.
Specifically, the selection is an iterative process consisting of $K$ steps. At each step, AVP proactively selects the most informative and not-yet-chosen region from the candidate set $C$ with the explicit goal of minimizing the uncertainty of the model’s answer to the question. The chosen region is then added to an intermediate set $R^{*}$. Notably, a key merit of AVP is its ability to bypass regions that, despite high attention scores, exhibit strong informational overlap with regions that have been selected. After $K$ rounds, the collection of all selected regions in $R^{*}$ constitutes the final optimal selection $S$. Substantial results in Appendices J.2 and L suggest that our framework’s average inference time is no more than 1.36 times that of an efficient baseline ICoT, and the framework scales efficiently with larger values of $K$ and $N_{C}=|C|$, respectively.
4.4 Dynamic Attention-shifting Trigger (DAT)
As motivated by the key observation in Section 3.3, it is vital to appropriately time the insertion of visual information when constructing a multimodal CoT, and furthermore, the shift of attention from the textual to the visual context serves as a crucial indicator. Motivated by these insights, we propose the DAT mechanism, which systematically evaluates the model’s attention scores on the visual context at every token $t$ generation step formulated as follows:
$A_{visual}(t)=\sum_{i\in\text{indices of $C_{visual}$}}\bar{a}_{t,i},$ (10)
where $C_{visual}$ is the visual information within the context; $\bar{a}_{t,i}$ is the average attention score of token $t$ towards the visual token with index $i$ across the last $N_{L}$ VLM layers. Drawing inspiration from the NLP community Jawahar et al. (2019); Tenney et al. (2019); Vig & Belinkov (2019), we restrict our focus to the model’s final layers, as they are responsible for capturing high-level semantic information, including semantic roles and coreference relations, where the shifting signal is presumed to be more reliable. In our implementation, we use the last 3 layers by default ($N_{L}=3$). Then, the shift of attention is formalized as follows:
$\Delta A_{visual}(t)=A_{visual}(t)-A_{visual}(t-1),$ (11)
which quantifies the model’s attention shift towards the visual context between generating the current token and the preceding one. Eventually, a hyper-parameter $\delta\in\mathds{R}$ is employed to delineate the point at which the attention shift $\Delta A_{visual}(t)$ is substantial enough to activate the AVP to insert essential visual information in the multimodal CoT. We detail a sensitivity analysis of the threshold $\delta$ with an emphasis on its impact on the frequency of triggering AVP and the performance of AIMCoT in Appendix K.
5 Experiments
5.1 Benchmarks and Baselines
In this study, we evaluate AIMCoT on three popular and challenging VQA benchmarks, including M3CoT Chen et al. (2024), ScienceQA Saikh et al. (2022), and LLaVA-Bench In-the-Wild (LLaVA-W) Liu et al. (2024). We provide detailed introductions in Appendix C.
To evaluate the performance of AIMCoT, we introduce the vanilla VLM w/o CoT (No-CoT) and a range of state-of-the-art methods as baseline models, including DDCoT Zheng et al. (2023), MMCoT Zhang et al. (2023), CCoT Mitra et al. (2024), and SCAFFOLD Lei et al. (2024), which generate text-only rationales. Furthermore, ICoT Gao et al. (2025), which constructs interleaved-modal CoT, is considered as well. The detailed introduction to them is listed in Appendix D. In presenting the results, we directly cite the performance reported in existing works where applicable.
5.2 Implementation Details
We implement AIMCoT and the baselines on Chameleon-7B Team (2024) and Qwen2-VL-7B-Instruct Wang et al. (2024) in two settings (both 0- and 1-shot), which aligns with the recent leading research Gao et al. (2025). The experiments are conducted on A6000 GPUs. The hyper-parameter settings for AIMCoT and the reproducibility statement are meticulously listed in Appendix G.2 and Section 7, respectively.
5.3 Performance Comparison
We evaluate the performance of AIMCoT against the state-of-the-art (SOTA) methods. The results shown in Table 2 clearly demonstrate the superiority of our proposed AIMCoT as it significantly outperforms all the baseline models under both 0- and 1-shot settings across all the datasets.
Specifically, AIMCoT surpasses all baselines that generate text-only rationales, confirming the efficacy of integrating salient visual information directly into CoT. When compared to ICoT, which also produces interleaved text-vision CoT, AIMCoT’s superior performance underscores the importance of our three key contributions: (1) a more reliable attention map as a foundation, (2) a proactive, goal-oriented mechanism for image region selection, and (3) an intelligent trigger for inserting visual information at critical moments.
Crucially, AIMCoT’s advantage is most pronounced on the open-ended LLaVA-W benchmark and in the 0-shot setting, which better simulate complex, real-world scenarios where the model must rely solely on its internal knowledge and reasoning. By emulating what can be seen as a more human-like cognitive process, AIMCoT unlocks the VLM’s foundational reasoning capabilities, enabling robust performance in novel and challenging situations.
@@ 5.3 Performance Comparison
Table 2: Performance comparison results on three widely-used benchmarks. The best performances are shown in bold. The metric for experiments on M3CoT and ScienceQA is Accuracy (ACC.), while on LLaVA-W, the metric ROUGE-L is adopted.
Backbone Method M3CoT (ACC.) ScienceQA (ACC.) LLaVA-W (ROUGE-L)
@@
AIMCoT (Ours) 31.4 32.8 53.1 54.5 29.8 32.0
Improvement 5.50% 1.47% 4.08% 2.04% 18.25% 15.94%
Qwen2-VL-7B No-CoT 43.6 45.4 56.3 64.4 32.7 33.5
@@
AIMCoT (Ours) 44.7 46.6 57.4 66.3 36.3 37.3
Improvement 1.4% 1.3% 1.1% 1.3% 6.2% 4.5%
5.4 Ablation Study
In this section, we conduct a series of ablation studies to verify the efficacy of each component within AIMCoT. The details of settings are as follows:
  • In w/o CAG, the VLM is directly prompted with the raw question $x$ and the paired image $I$;
  • In w/o AVP, the AVP is replaced by the attention-driven Top-K strategy by following existing works Gao et al. (2025); Ge et al. (2025), which selects the regions with Top-K attention scores on the model’s attention map;
  • In w/o DAT, following existing research Gao et al. (2025), the insertion of visual information in CoT is triggered when the model outputs the signal token, which is a line break by default.
Ablation results in Table 3 validate the contributions of our core components. First, CAG provides essential context enhancement, proving crucial for generating high-quality CoT, particularly when text queries are sparse (M3CoT, LLaVA-W). Second, the consistent, significant performance drop when replacing AVP with a Top-K baseline underscores the substantial superiority of our proactive, information-oriented method. Furthermore, the removal of DAT also results in a considerable performance decline, highlighting the critical importance of the timing of image insertion.
Table 3: Ablation study of AIMCoT conducted on Chameleon-7B under 0-shot setting.
Dataset AIMCoT w/o CAG w/o AVP w/o DAT

### v2

@@ 4.1 Preliminaries
Due to space constraints, we defer the definitions of important concepts to Appendix B, including (1) the Vision-Language Models (VLMs), (2) the Context Window of a VLM, (3) Patches and Regions (i.e., Visual Tokenization), and (4) I-MCoT.
@@ 4.2 Context-enhanced Attention-map Generation (CAG)
As revealed in Section 3.1, attention is an unreliable basis for visual selection, particularly under severe text-vision granularity imbalance. Although AIM-CoT shifts from passive attention-based selection to proactive information-driven selection, we argue that refining the VLM’s attention distribution remains meaningful for the interleaved-modal reasoning process. This is primarily for two reasons: (1) a more reliable attention distribution benefits reasoning, since it serves as the basis of model generation; and (2) crucially, the attention map serves as a source of candidate regions for subsequent selection (Section 4.3).
To this end, we propose CAG to improve the VLM’s cross-attention distribution by mitigating the text-vision granularity imbalance. Specifically, before the VQA process begins, the VLM is elicited to carefully generate an explanatory description of the input image conditioned on the textual query. Rather than appending a caption, this process enriches the brief query with more semantic anchors that encode visual information within the image, thereby alleviating the granularity imbalance. Formally, this is expressed as follows:
$\mathcal{D}_{\text{CAG}}$ $=\text{VLM}(I,x,\mathcal{P}_{\text{CAG}}),$ (1)
$x^{\prime}$ $=\text{concat}(x,\mathcal{D}_{\text{CAG}}),$ (2)
where $\mathcal{P}_{\text{CAG}}$ is the prompt for generating the description. $x^{\prime}$ is the updated textual query that integrates the description $\mathcal{D}_{\text{CAG}}$. A template of $\mathcal{P}_{\text{CAG}}$ is provided in Appendix A.3.
To ensure the output $\mathcal{D}_{\text{CAG}}$ serves as a reliable textual anchor while preventing error propagation to downstream modules, it is vital to mitigate potential hallucinations within it. To this end, the prompt $\mathcal{P}_{\text{CAG}}$ is meticulously designed with negative constraints: it explicitly instructs the model to prioritize visible evidence and adopt a rigorous, cautious stance—skipping uncertain details rather than speculating. Experiments in Appendix L.1 validate the effectiveness of negative constraints.
@@ 4.3 Active Visual Probing (AVP)
Drawing on the analyses in Section 3.2, we propose AVP, an information-driven mechanism designed to proactively select the regions with valuable information. Inspired by Information Foraging Theory (IFT), AVP operates through three systematic steps: (1) constructing a diverse candidate region set, (2) quantifying the information gain of each candidate region, and (3) executing a sequential, greedy selection process.
Step 1: Candidate Region Set Construction. As introduced in Appendix B, each input image is partitioned into multiple regions with a fixed partitioning method. Although the partition for a given input image is deterministic, evaluating the information gain of each of these regions could be computationally prohibitive. Therefore, it is necessary to construct a candidate pool, which is a subset of the raw partition.
We propose that although the candidate region set is a subset of the raw partition, it should remain sufficiently informative for subsequent selection. To this end, we derive candidate regions from diverse sources, covering both the VLM’s current focus and potential blind spots, as follows:
  • Attention-driven Set ($C_{attn}$): We select the top-$N$ regions from the CAG-enhanced cross-attention map ($A^{\prime}$), as this subset is a real-time reflection of VLM’s internal state in the latest reasoning step. Although the attention-based signals can inherently carry noise, analysis in Appendix I.3 suggests that our subsequent information-driven selection effectively filters out non-informative high-attention regions (i.e., noise rejection).
  • Exploratory Set ($C_{exp}$): To mitigate the “tunnel vision” of attention maps, we extract $M$ exploratory regions by uniformly sampling from all the partitioned regions. Despite its simplicity, this effective strategy outperforms complex alternatives (Appendix I.1) and provides essential regions overlooked by the attention-driven set $C_{attn}$ (Appendix I.2).
Formally, the total candidate set $C$ is constructed as follows:
$C_{attn}$ $=\{R_{1},R_{2},\dots,R_{N}\},$
$C_{exp}$ $=\{R_{N+1},R_{N+2},\dots,R_{N+M}\},$
$C$ $=C_{attn}\cup C_{exp},$
where each item denotes a specific image sub-region. For $i\leq N$, $R_{i}$ is the $i$-th most attended region in $A^{\prime}$.
Step 2: Information Gain Quantification. As analyzed in Section 3.2, IFT motivates us to identify the most informative regions. Therefore, it is necessary to quantify the information gain of each candidate region in $C$.
To quantify Information Gain (IG) of a candidate region, we compare the VLM’s information content before and after integrating this specific candidate (as visual tokens) into its current context. These pre- and post-insertion information contents are termed basic uncertainty and conditional uncertainty, respectively. Intuitively, they measure the VLM’s uncertainty in predicting the next token (textual) during autoregression under these two states, respectively. Formally, Basic Uncertainty ($U_{B}$) is defined as the entropy of the VLM’s next-token distribution given the current context, and Conditional Uncertainty ($U_{C,i}$) as the entropy after explicitly introducing candidate region $R_{i}$.
$U_{B}=H(Y|I,x,y_{<t})$
$=-\sum_{y\in V}P(y|I,x,y_{<t})\log_{2}P(y|I,x,y_{<t}),$ (3)
$U_{C,i}=H(Y|I,x,y_{<t},R_{i})$
$=-\sum_{y\in V}P(y|I,x,y_{<t},R_{i})\log_{2}P(y|I,x,y_{<t},R_{i}),$
where $V$ is the vocabulary, and $y_{<t}$ represents the generated tokens. The information gain is thus derived as:
$IG(\{R_{i}\})=U_{B}-U_{C,i},\;i=1,\dots,N+M.$ (4)
Although this calculation requires a forward pass, Appendix M shows that AIM-CoT remains highly deployable, with an average inference time under $1.36\times$ that of the highly efficient baseline.
Step 3: Sequential Selection Process. Finally, as analyzed in Section 3.2, drawing on the iterative nature of information foraging, where each step depends on previous knowledge, we frame region selection as a sequential trajectory. Specifically, in each step, we propose to select the region that provides the maximum information gain conditioned on the context updated in the last step, as outlined in Algorithm 1. This is because this greedy algorithm (1) intuitively moves maximally toward providing the VLM with the most visual information at each step, and besides, (2) it has theoretical support: regarding such a subset selection problem, it yields an approximation to the global optimum (i.e., the maximum information gain) for the VLM. We provide comprehensive analyses of this theoretical property in Appendix I.4.
1
Input : Total candidate set $C$, Target selection size $K$
Output : Optimal selection set $S$
2
[…]
4 for $k\leftarrow 1$ to $K$ do
5   Compute Basic Uncertainty: $U_{B}\leftarrow H(Y|I,x,y_{<t},R^{*})$ ;
6 for $R_{i}\in C\setminus R^{*}$ do
7      Compute Conditional Uncertainty: $U_{C,i}\leftarrow H(Y|I,x,y_{<t},R^{*}\cup\{R_{i}\})$ ;
8      Calculate Gain: $IG(\{R_{i}\})\leftarrow U_{B}-U_{C,i}$ ;
9
10  Select Best Region: $R_{next}\leftarrow\operatorname{argmax}_{R_{i}\in C\setminus R^{*}}\{IG(\{R_{i}\})\}$ ;
11   Update Selected Set: $R^{*}\leftarrow R^{*}\cup\{R_{next}\}$ ;
12
[…]
return $S$
Algorithm 1 Greedy Algorithm for Optimal Region Selection
4.4 Dynamic Attention-shift Trigger (DAT)
As identified in Section 3.3, although attention remains an unreliable basis for selection, its shift serves as a reliable diagnostic signal for triggering. Motivated by this, we propose DAT, the trigger component of AIM-CoT. DAT tracks the VLM’s text-to-vision attention shifts across the autoregressive steps, as they reflect the VLM’s dynamic need for visual evidence. Then, based on the magnitude of each shift, DAT determines whether the visual evidence should be introduced via AVP at that step.
Quantification of Text-to-vision Attention Shift.
The text-to-vision attention shift is determined by the difference in attention distributions between steps. Intuitively, it tracks the flow of attention from the text context in the previous step to the visual context in the current step. Furthermore, since the sum of attention scores over these two parts (i.e., text and vision) is fixed in the VLM’s normalized attention map, it is sufficient to focus on the attention shift in the visual part. Formally, this can be quantified as follows:
$\Delta A_{vision}(t)=A_{vision}(t)-A_{vision}(t-1),$ (5)
where $A_{vision}(t)$ and $A_{vision}(t-1)$ represent the sum of attention scores allocated to all the visual tokens in the context by the currently predicted token (the $t$-th token) and the preceding token (the (t1)(t-1)-th token), respectively.
Triggering Condition.
DAT determines if a given shift $\Delta A_{vision}(t)$ is significant enough to warrant the provision of visual evidence via a qualitative criterion. In the criterion, a hyper-parameter $\delta\in\mathbb{R}$ is introduced to serve as the threshold. The AVP module is triggered to perform visual insertion after the $t$-th generation step if and only if $\Delta A_{vision}(t)>\delta$.
In Appendix G.1, we provide a detailed sensitivity analysis of $\delta$ and introduce an adaptive thresholding strategy to set $\delta$ automatically.
Critical Integration Strategy.
A potential challenge inherent to the I-MCoT paradigm is the difficulty in eliminating noise that visual evidence might introduce. To address this, DAT does not naively trigger the visual evidence insertion. Instead, it employs a critical integration mechanism: the inserted regions are accompanied by a safety instruction, which cues the VLM to treat the visual evidence as “supplementary references”, encouraging the model to verify semantic consistency with the existing textual context. This effectively filters out irrelevant or hallucinated visual cues.
We validate the effectiveness of this safety instruction in Appendix L.2.
@@ 4.4 Dynamic Attention-shift Trigger (DAT)
Table 1: Performance comparison results on three VQA benchmarks and four backbones. The best performances are shown in bold. We report Accuracy (ACC.) for M3CoT and ScienceQA, and ROUGE-L for LLaVA-W.
Method M3CoT ScienceQA LLaVA-W
@@
AIM-CoT (ours) 31.4 32.8 53.1 54.5 29.8 32.0
Improv. +5.4% +1.5% +4.1% +2.1% +18.3% +15.9%
[…]
No-CoT 36.5 37.8 53.9 61.2 29.8 30.9
DDCoT 36.8 38.6 53.2 61.6 29.0 30.4
MMCoT 35.2 36.6 50.8 58.0 28.2 28.9
CCoT 36.4 38.2 54.1 61.5 27.8 31.5
SCAFFOLD 36.1 38.0 52.4 60.5 29.7 30.8
ICoT 37.6 39.4 55.1 62.5 32.5 33.6
AIM-CoT (ours) 39.7 41.5 56.9 64.9 35.5 36.6
Improv. +5.6% +5.3% +3.3% +3.8% +9.2% +8.9%
Method M3CoT ScienceQA LLaVA-W
0-shot 1-shot 0-shot 1-shot 0-shot 1-shot
[…]
No-CoT 43.6 45.4 56.3 64.4 32.7 33.5
DDCoT 42.6 45.7 55.2 64.9 31.2 32.8
@@
AIM-CoT (ours) 44.7 46.6 57.4 66.3 36.3 37.3
Improv. +1.4% +1.3% +1.1% +1.4% +6.1% +4.5%
[…]
No-CoT 55.2 56.8 73.5 77.2 40.5 41.8
DDCoT 54.8 57.5 72.9 77.8 39.6 41.3
MMCoT 53.4 54.7 69.8 74.5 38.9 39.2
CCoT 55.3 56.9 73.4 77.1 37.8 42.4
SCAFFOLD 54.2 56.7 71.6 76.3 40.1 41.5
ICoT 56.9 59.1 75.1 79.2 43.4 44.7
AIM-CoT (ours) 58.7 61.2 76.8 81.3 46.5 49.1
Improv. +3.2% +3.6% +2.3% +2.7% +7.1% +9.8%
Statistical significance: $^{\dagger}p<0.05$, $^{\ddagger}p<0.01$ for AIM-CoT compared with the second-best method under the same backbone/setting (McNemar for ACC. on M3CoT/ScienceQA; Wilcoxon signed-rank for ROUGE-L on LLaVA-W). Full results are reported in Appendix E.
5 Experiments
5.1 Evaluation Setup
Benchmark. We evaluate AIM-CoT on three widely used VQA benchmarks: M3CoT Chen et al. (2024a), ScienceQA Saikh et al. (2022), and LLaVA-W Liu et al. (2024). Detailed descriptions are provided in Appendix A.1.
Baselines. We compare against several text-only baselines, including vanilla VLM without CoT (No-CoT), DDCoT Zheng et al. (2023), MMCoT Zhang et al. (2024), CCoT Mitra et al. (2024), and SCAFFOLD Lei et al. (2025). On top of these, the leading I-MCoT framework ICoT Gao et al. (2025), is also included. More details are available in Appendix A.2. When possible, we directly report figures from prior work.
Backbones. The implementation is built on four mainstream VLM backbones with different architectures and scales: early-fusion Chameleon-7B Team (2024) and Janus-Pro-7B Chen et al. (2025), and late-fusion Qwen2-VL-7B Wang et al. (2024) and Qwen2.5-VL-32B Bai et al. (2025). We conduct experiments in both 0- and 1-shot settings, using the prompt template from the open-source ICoT implementation Gao et al. (2025).
Hyper-parameter. Hyper-parameter settings are listed in Appendix A.4.
Table 2: Ablation study of AIM-CoT conducted on Chameleon-7B under 0-shot setting.
Dataset AIM-CoT w/o CAG w/o AVP w/o DAT